OpenStudy (anonymous):

Given the trinomial 5x2 - 2x - 3, predict the type of solutions. Two rational solutions One rational solution Two irrational solutions Two complex solutions I don't understand!

hartnn (hartnn):

for a trinomial \(\Large ax^2 +bx+c\) find out \(\huge D = b^2 -4ac\)

OpenStudy (anonymous):

okay what would d stand for though?

hartnn (hartnn):

if D is Negative, then 2 complex solutions if D is Positive, then Real Solutions if D is a perfect square , then Rational Solutions IF D is =0 , then only one solution

hartnn (hartnn):

compare \(\Large 5x^2 -2x-3\) with \(\Large ax^2+bx+c\) fins a,b,c can you ?

OpenStudy (anonymous):

okay

hartnn (hartnn):

**find a,b,c

hartnn (hartnn):

need help with that ?

OpenStudy (anonymous):

no sorry for taking so long i was talking to someone

OpenStudy (anonymous):

a=5 b=-2 c=-3 Is this correct?

hartnn (hartnn):

absolutely correct :)

OpenStudy (anonymous):

cool okay let me try this

hartnn (hartnn):

now find \(D = b^2-4ac = ..\)

OpenStudy (anonymous):

\[D=-2^{2}-4\left( 5 \right)\left( -3 \right)\]

OpenStudy (anonymous):

\[D=-2^{2}-4\left( -15 \right)\]

hartnn (hartnn):

careful with the brackets ! \(D = (-2)^2 - 4 (5)(-3)\)

OpenStudy (anonymous):

\[D=-2^{2}+50\]

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

\[D=\left( -2 \right)^{2}+50\]

OpenStudy (anonymous):

\[D=-4+50\]

hartnn (hartnn):

noooo

OpenStudy (anonymous):

oh crap

hartnn (hartnn):

its \(\Large b^2\) and b = -2 so it'll be \((-2)^2 \)

hartnn (hartnn):

\((-2)^2 = (-2)\times (-2) =+4\) makes sense ?

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so i just had to make it positive

OpenStudy (anonymous):

\[D=4+50=54\]

hartnn (hartnn):

\(\Large -4 (5)(-3) = -4 (-15) = +60\)

OpenStudy (anonymous):

wow i was really off

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

okay so it would be 64

hartnn (hartnn):

not much :) mistakes happen, that only means you're learning :)

hartnn (hartnn):

yes D =64

OpenStudy (anonymous):

so it would be b?

hartnn (hartnn):

now what is D (64) among these? if D is Negative, then 2 complex solutions if D is Positive, then Real Solutions if D is a perfect square , then Rational Solutions IF D is =0 , then only one solution

OpenStudy (anonymous):

b

hartnn (hartnn):

b : One rational Solution ??? NO we have one solution when D =0

OpenStudy (anonymous):

but what does it mean if it is a perfect square?

hartnn (hartnn):

is 64 a perfect square ? or you wanna know what is "perfect square"

OpenStudy (anonymous):

what would be considered a perfect square?

hartnn (hartnn):

a perfect square is a number whose square root is an integer

hartnn (hartnn):

like 25 is a perfect square because \(\sqrt {25} = 5\)

hartnn (hartnn):

whether 2 is not a perfect square because \(\sqrt {2} = 1.414\) and 1.414 is not an integer

hartnn (hartnn):

whats \(\sqrt {64} =.. ?\) you know ?

OpenStudy (anonymous):

8

hartnn (hartnn):

is 8 an integer ? if yes, then 64 is a perfect square :)

OpenStudy (anonymous):

8 is an integer so you wouldnt have to completely break it down in order to know?

hartnn (hartnn):

i didn't get your question .. .

hartnn (hartnn):

if D is a perfect square , then 2 Rational Solutions

OpenStudy (anonymous):

oh i got it okay thank you so much for your help!!!

hartnn (hartnn):

welcome ^_^